# Neural network for excess noise estimation in continuous-variable quantum key distribution under composable finite-size security

### v2

@@ I Introduction
In a generic CV-QKD protocol, quantum information can be encoded onto coherent states by modulating the amplitude and phase quadratures of laser light, typically using electro-optical modulators at the transmitter to establish a secret key between two legitimate QKD users (Alice and Bob) Grosshans and Grangier (2002a); Grosshans et al. (2003a); Weedbrook et al. (2004). These states are transmitted through a quantum channel that is assumed to be fully under the control of a potential eavesdropper, conventionally referred to as Eve. The security of CV-QKD protocols employing Gaussian-modulated coherent states was initially proven in the asymptotic limit Van Assche et al. (2005); Navascués et al. (2006); García-Patrón and Cerf (2006) and later extended to the finite-size regime, ensuring universal composability against both collective Leverrier et al. (2010) and general coherent attacks Furrer et al. (2012). Beyond continuous modulation, discrete modulation of coherent states has also been extensively studied and shown to offer promising performance and security guarantees Leverrier and Grangier (2009); Denys et al. (2021); Bäuml et al. (2024).
@@ II Protocol and model definition
In this work, we will consider the coherent-state protocol with Gaussian modulation Grosshans and Grangier (2002b). The protocol starts with Alice preparing $N$ displaced vacuum states $\lvert q_{i}+ip_{i}\rangle$ by modulating both the amplitude and phase quadratures. The displacements $q_{i}$ and $p_{i}$ are independent random variables drawn from the normal distribution $\mathcal{N}(0,V_{A})$. These states are transmitted over an untrusted channel with transmission $T$ and excess noise $\xi$, assumed to be under Eve’s control. Upon reception, Bob performs homodyne detection by switching randomly between the phase and amplitude quadratures. Here, Alice’s use of modulation to prepare these states constitutes a prepare-and-measure protocol, known to have an entanglement-based equivalent Grosshans et al. (2003b), which is used for the security analysis García-Patrón and Cerf (2006).
@@
$T\approx\eta_{\text{eff}}10^{-0.02d}$ (2)
where $\eta_{\text{eff}}$ is the known quantum efficiency of Bob’s detection and $d$ is the distance in kilometers between Alice and Bob. In this case, we assume an optical fiber with losses of $0.2$ dB per kilometer. More generally, we consider the possibility that Eve can access side-channel information resulting from imperfections in the detection setup. In this analysis, we focus on the receiver’s detection inefficiency and assume that the fraction $1-\eta_{\text{eff}}$ of the incoming photons that are not detected is not simply lost to the environment, but is instead collected by Eve and incorporated into her attack strategy Pirandola (2021).
@@
where $I(x:y)$ denotes the mutual information between Alice’s and Bob’s classical variables $x$ and $y$ Cover (1999), while $\chi(y:E)$ represents the Holevo information between Bob’s variable $y$ and the adversary’s quantum system $E$ Nielsen and Chuang (2010). The supremum is taken over all channels $\mathcal{N}:A\rightarrow B$ that are consistent with the statistics observed by Alice and Bob during the parameter estimation. For the entanglement-based protocol Grosshans et al. (2003b), it can be described by the covariance matrix
@@
where $d$ denotes the number of bits per quadrature used during discretization, $\bar{\epsilon}$ is the smoothing parameter, and $\epsilon_{\mathrm{PA}}$ is the failure probability of the privacy‐amplification procedure Leverrier et al. (2010). Both $\bar{\epsilon}$ and $\epsilon_{\mathrm{PA}}$ are intermediate quantities to be optimized numerically. The first term of $\Delta(n)$, namely the square‐root term, quantifies the convergence rate of the smooth min‐entropy — the relevant metric for key length — of an i.i.d. state under collective attacks toward its von Neumann entropy, since only in the asymptotic limit does the smooth min‐entropy equal the von Neumann entropy. Its derivation was done in ref. Pirandola (2021), which used the framework proposed in ref. Tomamichel (2013). The second term directly reflects the security contribution of the failure probability $\epsilon_{\mathrm{PA}}$ in the privacy‐amplification step.
@@ III Parameter estimation effects on protocol security
The main problem here is that Bob and Alice do not know these parameters, since it is assumed that the channel can be freely controlled by Eve. The law of large numbers guarantees that when $m\rightarrow\infty$ $\mathbb{E}[\hat{t}]\equiv t$ and $\mathbb{E}[\hat{\sigma}^{2}]\equiv\sigma^{2}$ Casella and Berger (2024), so there is no need for error analysis in asymptotic limit. In practical implementations, it is obviously impossible to achieve this result, such that one needs to consider the probability of failure of the parameter estimation $\epsilon_{PE}$ using statistical analysis for the estimators.
The maximum likelihood estimation (MLE) is widely recognized as the standard method in the field, since it is compatible with statistical analyses considering the confidence interval Thearle et al. (2016); Pirandola and Papanastasiou (2024); Papanastasiou et al. (2017); Zhang et al. (2020). For the linear model,
@@
$\sigma^{2}_{\max-\text{MLE}}\approx\hat{\sigma}^{2}_{MLE}+z_{\epsilon_{PE}/2}\frac{\hat{\sigma}^{2}\sqrt{2}}{\sqrt{m}}\,.$ (10)
where $z_{\epsilon_{PE}/2}=\text{erf}^{-1}(1-\epsilon_{PE}/2)$ and $\text{erf}(x)$ is the error function.
@@
which means that there exists a confidence set $\mathcal{C}_{\epsilon_{PE}}$ such that the covariance matrix $\Gamma_{\epsilon_{PE}}$ lies within $\mathcal{C}_{\epsilon_{PE}}$ with probability at least $1-\epsilon_{PE}/2$. Thus, the secret key rate that accounts for the probability of failure in parameter estimation can be computed using PEP 1.
Parameter estimation protocol 1
- Parameter estimation via maximum likelihood estimation in the finite-size scenario
  1. 1.
    Since Alice only has access to his measured signals $y$, Bob needs to broadcast $m$ signals over an authenticated channel so that Bob can estimate $t$ and $\sigma^{2}$.
  2. 2.
    Alice uses estimator from Eq. (8) to estimate $t$ and $\sigma^{2}$, using the $m$ correlated data.
  3. 3.
    Alice uses the statistical analysis from Eq. (9) to compute $t_{min}$ and from Eq. (10) to compute $\sigma^{2}_{max}$.
  4. 4.
    Alice uses these results to write the covariance matrix from Eq. (11) and, finally, compute $\chi_{\epsilon_{PE}}(y:E)$.
@@ IV Worst-case confidence intervals for neural networks
where $\mathrm{X}_{i}$ denotes the input corresponding to the $i$-th observation Seber and Wild (2005). This modeling framework is well established in computational learning theory, as it enables statistical generalizations across a broad class of inference tasks Barron (1994); Papadopoulos et al. (2001). The central challenge, therefore, lies in demonstrating that neural networks can be effectively described within this framework under appropriate assumptions, thereby ensuring reliable and secure parameter estimation. In this work, we adopt such a perspective, drawing inspiration from the delta method outlined in refs. Chryssolouris et al. (1996); Hwang and Ding (1997).
@@
This procedure is expected to bring $\hat{\theta}$ closer to $\theta^{*}$. In principle, minimizing the loss function corresponds to finding the optimal set of network parameters that best approximates the underlying functional relationship between the inputs and the outputs Li et al. (2018). Given a sufficiently expressive architecture and representative training data, the neural network learns a map that minimizes the discrepancy between the predicted and true values of the target variable.
@@
The first term corresponds to the intrinsic measurement noise, while the second term captures the uncertainty in the model prediction due to the estimation error in the parameters. The minimization process justifies modeling the noise $\varepsilon_{0}$ as a zero-mean Gaussian random variable with variance $\sigma^{2}_{\varepsilon}$, such that the parameter estimation error $\hat{\theta}-\theta^{*}$ can be approximated as following a multivariate normal distribution
@@
which is constant for a given $\mathrm{Y}\in\mathcal{S}$, where $S$ delimits the set of possible values for the excess noise estimation. This variance estimator is then used in the Student’s $t$-distribution to quantify the uncertainty associated with the prediction $\hat{\mathrm{Y}}_{0}$. Thus, one can construct a confidence interval for the predicted value $\hat{\mathrm{Y}}_{0}$, which incorporates both the estimated variance and the sensitivity of the prediction to the parameters through the Jacobian vector $\mathbf{f}_{0}$:
@@
From a computational picture, the trained neural network is defined as a family of probability distributions on a sample space of excess noise $S$, indexed by a parameter vector $\hat{\theta}\in\Theta$. Thus, the trained neural network acts as a statistical estimator $\hat{\theta}:S\to\Theta$, approximating the mapping from data to channel parameters. The quantity in Eq. (24) defines a confidence interval with probability $1-\epsilon/2$, which is operationally equivalent to the parameter estimation method based on MLE in ref. Leverrier et al. (2010):
@@
Since neural networks can be computationally expensive, the efforts invested in them must be used on processes that have significant impacts on the key rate. As discussed, this is the case for excess noise Huang et al. (2016); Laudenbach et al. (2018); Hajomer et al. (2024). In this case, the neural network to estimate the variance in the worst-case is then defined, and finally, we can estimate the excess noise via PEP 2, which is operationally equivalent to PEP 1.
Parameter estimation protocol 2
- Parameter estimation via neural network in the finite-size scenario
  1. 1.
    Since Alice only has access to his measured signals $y$, Bob needs to broadcast $m$ signals over an authenticated channel so that Bob can estimate $t$ and $\sigma^{2}$.
  2. 2.
    Alice uses estimator from Eq. (8) to compute $\hat{t}$ and a trained neural network to compute $\hat{\sigma}^{2}$.
  3. 3.
    Alice uses the statistical analysis from Eq. (9) to compute $t_{min}$ and from Eq. (24) to compute $\sigma^{2}_{max}$.
  4. 4.
    Alice uses these results to write the covariance matrix from Eq. (11) and, finally, compute $\chi_{\epsilon_{\mathrm{PE}}}(y:E)$.
@@ V Neural network model
In this section, we present the neural network architecture developed to implement the framework introduced previously. We detail the data input structure, the network’s architectural design, and the training strategy employed. A discussion regarding the computational cost associated with the proposed neural network is provided in Appendix A.
@@ V.1 Neural network inputs
To estimate the noise variance parameter in a CV-QKD system, we designed a fully-connected feedforward neural network tailored to extract nonlinear correlations from statistical quadrature measurements of the channel. All input data are expressed in shot-noise units (SNU), and the network is trained to estimate the product $\hat{t}^{2}\hat{\xi}$, since the parameter $\mu$ is constant for diferent frames.
@@
$\text{Var}(y^{\prime})=(\hat{t}-\hat{t}_{\text{MLE}})^{2}V_{A}+\frac{\hat{t}_{\text{MLE}}^{2}}{a^{2}}V_{A}+\hat{t}^{2}\xi+1,$ (27)
@@
$a^{2}(\text{Var}(y^{\prime})-1)\approx\hat{t}_{\text{MLE}}^{2}V_{A}+a^{2}\hat{t}^{2}\xi,$ (28)
since the discrepancy $(\hat{t}-\hat{t}_{\text{MLE}})^{2}$ vanishes in the large-sample limit due to the consistency of the MLE.
The covariance term $\text{Cov}(x,y^{\prime})$ is computed from the sample using the standard Pearson correlation estimator. The output of the network is then post-processed by dividing by $a^{2}$, recovering an accurate estimate of the original parameter $\hat{t}^{2}\hat{\xi}$ from the amplified noise features.

### v3

@@ I Introduction
In a generic CV-QKD protocol, quantum information can be encoded onto coherent states by modulating the amplitude and phase quadratures of laser light, typically using electro-optical modulators at the transmitter to establish a secret key between two legitimate QKD users (Alice and Bob) Grosshans and Grangier (2002b); Grosshans et al. (2003b); Weedbrook et al. (2004). These states are transmitted through a quantum channel that is assumed to be fully under the control of a potential eavesdropper, conventionally referred to as Eve. The security of CV-QKD protocols employing Gaussian-modulated coherent states was initially proven in the asymptotic limit Van Assche et al. (2005); Navascués et al. (2006); García-Patrón and Cerf (2006) and later extended to the finite-size regime, ensuring universal composability against both collective Leverrier et al. (2010) and general coherent attacks Furrer et al. (2012). Beyond continuous modulation, discrete modulation of coherent states has also been extensively studied and shown to offer promising performance and security guarantees Leverrier and Grangier (2009); Denys et al. (2021); Bäuml et al. (2024).
@@ II Protocol and model definition
In this work, we will consider the coherent-state protocol with Gaussian modulation Grosshans and Grangier (2002a). The protocol starts with Alice preparing $N$ displaced vacuum states $\lvert q_{i}+ip_{i}\rangle$ by modulating both the amplitude and phase quadratures. The displacements $q_{i}$ and $p_{i}$ are independent random variables drawn from the normal distribution $\mathcal{N}(0,V_{A})$. These states are transmitted over an untrusted channel with transmission $T$ and excess noise $\xi$, assumed to be under Eve’s control. Upon reception, Bob performs homodyne detection by switching randomly between the phase and amplitude quadratures. Here, Alice’s use of modulation to prepare these states constitutes a prepare-and-measure protocol, known to have an entanglement-based equivalent Grosshans et al. (2003a), which is used for the security analysis García-Patrón and Cerf (2006).
@@
$T\approx\eta_{\text{eff}}10^{-0.02\ell}$ (2)
where $\eta_{\text{eff}}$ is the known quantum efficiency of Bob’s detection and $\ell$ is the distance in kilometers between Alice and Bob. In this case, we assume an optical fiber with losses of $0.2$ dB per kilometer. More generally, we consider the possibility that Eve can access side-channel information resulting from imperfections in the detection setup. In this analysis, we focus on the receiver’s detection inefficiency and assume that the fraction $1-\eta_{\text{eff}}$ of the incoming photons that are not detected is not simply lost to the environment, but is instead collected by Eve and incorporated into her attack strategy Pirandola (2021).
@@
where $I(x:y)$ denotes the mutual information between Alice’s and Bob’s classical variables $x$ and $y$ Cover (1999), while $\chi(y:E)$ represents the Holevo information between Bob’s variable $y$ and the adversary’s quantum system $E$ Nielsen and Chuang (2010). The supremum is taken over all channels $\mathcal{N}:A\rightarrow B$ that are consistent with the statistics observed by Alice and Bob during the parameter estimation. For the entanglement-based protocol Grosshans et al. (2003a), it can be described by the covariance matrix
@@
where $d$ denotes the number of bits per quadrature used during discretization (e.g., $d=2^{5}=32$ for a 5-bit discretization), $\bar{\epsilon}$ is the smoothing parameter, and $\epsilon_{\mathrm{PA}}$ is the failure probability of the privacy‐amplification procedure Leverrier et al. (2010). Both $\bar{\epsilon}$ and $\epsilon_{\mathrm{PA}}$ are intermediate quantities to be optimized numerically. The first term of $\Delta(n)$, namely the square‐root term, quantifies the convergence rate of the smooth min‐entropy — the relevant metric for key length — of an i.i.d. state under collective attacks toward its von Neumann entropy, since only in the asymptotic limit does the smooth min‐entropy equal the von Neumann entropy. Its derivation was done in ref. Pirandola (2021), which used the framework proposed in ref. Tomamichel (2013). The second term directly reflects the security contribution of the failure probability $\epsilon_{\mathrm{PA}}$ in the privacy‐amplification step.
@@ III Parameter estimation effects on protocol security
The main problem is that Bob and Alice do not know these parameters, since it is assumed that the channel can be freely controlled by Eve. The law of large numbers guarantees that when $m\rightarrow\infty$ $\mathbb{E}[\hat{t}]\equiv t$ and $\mathbb{E}[\hat{\sigma}^{2}]\equiv\sigma^{2}$ Casella and Berger (2024), so there is no need for error analysis in asymptotic limit. In practical implementations, it is obviously impossible to achieve this result, such that one needs to consider the probability of failure of the parameter estimation $\epsilon_{PE}$ using statistical analysis for the estimators.
The maximum likelihood estimation is widely recognized as the standard method in the field, since it is compatible with statistical analyses considering the confidence interval Thearle et al. (2016); Pirandola and Papanastasiou (2024); Papanastasiou et al. (2017); Zhang et al. (2020). For the linear model,
@@
$\sigma^{2}_{\max-\text{MLE}}\approx\hat{\sigma}^{2}_{MLE}+z_{\epsilon_{PE}/2}\frac{\hat{\sigma}^{2}\sqrt{2}}{\sqrt{m}}\,$, (10)
where $z_{\epsilon_{PE}/2}=\text{erf}^{-1}(1-\epsilon_{PE}/2)$ and $\text{erf}(x)$ is the error function, ensuring that the key is not overestimated by one of the authenticated parties with probability at least $1-\epsilon_{PE}/2$, i.e.,
$P(\sigma^{2}\leq\sigma_{\mathrm{max}}^{2})\geq 1-\epsilon_{\mathrm{PE}}/2\quad$ (11)
and
$P(t\geq t_{\mathrm{min}})\geq 1-\epsilon_{\mathrm{PE}}/2\,.$ (12)
@@
which means that there exists a confidence set $\mathcal{C}_{\epsilon_{PE}}$ such that the covariance matrix $\Gamma_{\epsilon_{PE}}$ lies within $\mathcal{C}_{\epsilon_{PE}}$ with probability at least $1-\epsilon_{PE}/2$. Thus, the secret key rate that accounts for the probability of failure in parameter estimation can be computed using PEP III.
PEP 1:  Parameter estimation via maximum likelihood estimation in the finite-size scenario 1. Since Alice only has access to her transmitted signals $y$, Bob needs to broadcast $m$ signals over an authenticated channel so that Alice can estimate $t$ and $\sigma^{2}$. 2. Alice uses estimator from Eq. (8) to estimate $t$ and $\sigma^{2}$, using the $m$ correlated data. 3. Alice uses the statistical analysis from Eq. (9) to compute $t_{min}$ and from Eq. (10) to compute $\sigma^{2}_{max}$. 4. Alice uses these results to write the covariance matrix from Eq. (13) and, finally, compute $\chi_{\epsilon_{PE}}(y:E)$.
@@ IV Worst-case confidence intervals for neural networks
where $\mathrm{X}_{i}$ denotes the input corresponding to the $i$-th observation Seber and Wild (2005). This modeling framework is well established in computational learning theory, as it enables statistical generalizations across a broad class of inference tasks Barron (1994); Papadopoulos et al. (2001). While statistical estimators derives confidence intervals directly from the known sampling distribution of the physical parameter estimators Casella and Berger (2024), machine learning-based estimation includes the explicit error term $\varepsilon$, as it accounts for the approximation error between the learning-theoretic framework’s predictions and the true underlying function.
The central challenge, therefore, lies in demonstrating that neural networks can be effectively described within this framework under appropriate assumptions, thereby ensuring reliable and secure parameter estimation. In this work, we adopt such a perspective, drawing inspiration from the delta method outlined in refs. Chryssolouris et al. (1996); Hwang and Ding (1997).
@@
This procedure is expected to bring $\hat{\theta}$ closer to $\theta^{*}$, where $\theta^{*}$ denotes the optimal set of network parameters that minimizes the expected prediction error and $\hat{\theta}$ is the set obtained by minimizing the empirical loss. In principle, minimizing the loss function corresponds to finding the optimal set of network parameters that best approximates the underlying functional relationship between the inputs and the outputs Li et al. (2018). Given a sufficiently expressive architecture and representative training data, the neural network learns a map that minimizes the discrepancy between the predicted and true values of the target variable.
@@
The first term corresponds to the intrinsic measurement noise, while the second term captures the uncertainty in the model prediction due to the estimation error in the parameters. The first-order Taylor expansion in Eq. (17) introduces a linearization error of order $\|\hat{\theta}-\theta^{*}\|^{2}$. This error becomes negligible in the large-sample regime where $\hat{\theta}$ converges to $\theta^{*}$. Moreover, the confidence intervals derived via the delta method are conservative by construction, meaning they account for such higher-order uncertainties and preserve the composable security of the estimation scheme.
In this framework, we introduce the hypothesis that the prediction error $\varepsilon_{0}$ follows a zero-mean Gaussian distribution based on two complementary considerations: the physical channel model and the statistical properties of the trained neural network. First, because the network is trained on data generated from the normal linear model, the resulting prediction errors are expected to approximately inherit Gaussian characteristics. Second, minimization of the MSE loss yields parameter estimates $\hat{\theta}$ that converge to a Gaussian random vector around its optimal value $\theta^{*}$. This follows from the fact that the loss gradient is a sum of many independent terms, inducing asymptotic normality via the central limit theorem. Considering this approach, the parameter estimation error $\hat{\theta}-\theta^{*}$ can be approximated as following a multivariate normal distribution
@@
Once the neural network is trained, the variance of the additive error term $\sigma^{2}_{\varepsilon}$ can be computed during the test process using an unbiased estimator based on the residual sum of squares:
@@
which is constant for a given $\mathrm{Y}\in\mathcal{S}$, where $\mathcal{S}$ delimits the set of possible values for the excess noise estimation. Its value is computed from the same synthetic dataset used to evaluate the full Jacobian matrix $\mathbf{F}$, while the Jacobian vector $\mathbf{f}_{0}$ corresponds to the gradient at the the actual input sample used to estimate the channel parameters $\mathbf{X}_{0}$. This separation allows the Jacobian matrix to be pre‑computed, while the lightweight vector $\mathbf{f}_{0}$ is evaluated per frame, ensuring statistical rigor through error propagation while maintaining real‑time feasibility.
Then the variance estimator is used in the Student’s $t$-distribution to quantify the uncertainty associated with the prediction $\hat{\mathrm{Y}}_{0}$. Thus, one can construct a confidence interval for the predicted value $\hat{\mathrm{Y}}_{0}$, which incorporates both the estimated variance and the sensitivity of the prediction to the parameters through the Jacobian vector $\mathbf{f}_{0}$:
@@
From a computational picture, the trained neural network is defined as a family of probability distributions on a sample space of excess noise $\mathcal{S}$, indexed by a parameter vector $\hat{\theta}\in\Theta$. Thus, the trained neural network acts as a statistical estimator $\hat{\theta}:\mathcal{S}\to\Theta$, approximating the mapping from data to channel parameters. The quantity in Eq. (26) defines a confidence interval with probability $1-\epsilon/2$:
@@
Since neural networks can be computationally expensive, the efforts invested in them must be used on processes that have significant impacts on the key rate. As discussed, this is the case for excess noise Huang et al. (2016); Laudenbach et al. (2018); Hajomer et al. (2024). In this case, the neural network to estimate the variance in the worst-case is then defined, and finally, we can estimate the excess noise via PEP IV.
PEP 2:  Parameter estimation via neural network in the finite-size scenario 1. Since Alice only has access to her transmitted signals $y$, Bob needs to broadcast $m$ signals over an authenticated channel so that Alice can estimate $t$ and $\sigma^{2}$. 2. Alice uses estimator from Eq. (8) to compute $\hat{t}$ and a trained neural network to compute $\hat{\sigma}^{2}$. 3. Alice uses the statistical analysis from Eq. (9) to compute $t_{min}$ and from Eq. (26) to compute $\sigma^{2}_{max}$. 4. Alice uses these results to write the covariance matrix from Eq. (13) and, finally, compute $\chi_{\epsilon_{\mathrm{PE}}}(y:E)$.
The quantity in Eq. (26) defines a confidence interval with probability $1-\epsilon_{PE}/2$ used in PEP IV, which is operationally equivalent to MLE-based parameter estimation method from PEP III. This equivalence operates on two fundamental levels:
  • Security Guarantee: Both methods produce an upper bound $\sigma^{2}_{\text{max}}$ that satisfies the same probabilistic condition $P(\sigma^{2}\leq\sigma^{2}_{\text{max}})\geq 1-\epsilon_{PE}/2$, which is the direct requirement for composable security in the parameter estimation stage.
  • Protocol Integration: The value $\sigma^{2}_{\text{max}}$ generated by either method is inserted into the same worst-case covariance matrix (Eq.13) and used to compute the same Holevo information quantity $\chi_{\epsilon_{PE}}(y:E)$ in the secret key rate (Eq. 6). From the perspective of the security protocol, the operation is indistinguishable, as shown in Fig. 1.
Figure 1: Schematic representation of the parameter estimation step in a CV-QKD protocol. The two parallel branches illustrate the operational equivalence between the conventional MLE and the proposed neural-network estimator. In both branches, Alice and Bob use the same set of correlated quadrature measurements $\{x_{i},y_{i}\}_{m}$ to compute a worst-case upper bound $\sigma^{2}_{\max}$ on the noise variance and a lower bound $t_{\min\text{-MLE}}$ on the channel transmittance. These bounds are then used to construct the same worst-case covariance matrix $\Gamma_{\epsilon_{PE}}$, from which the smooth Holevo information $\chi_{\epsilon_{PE}}(y:E)$ entering the secret key rate is evaluated.
@@ V Neural network model
In this section, we present the neural network architecture developed to implement the framework introduced previously. We detail the data input structure, the network’s architectural design, and the training strategy employed. A discussion regarding the computational cost associated with the proposed neural network is provided in Appendix A, while the practical implementation for confidence interval is provided in Appendix B.
@@ V.1 Neural network inputs
To estimate the noise variance parameter in a CV-QKD system, we designed a fully-connected feedforward neural network tailored to extract nonlinear correlations from statistical quadrature measurements of the channel. All input data are expressed in shot-noise units (SNU), and the network is trained to estimate the product $\hat{t}^{2}\hat{\xi}$, since the parameter $\mu$ is constant for different frames.
@@
$\text{Var}(y^{\prime})=(\hat{t}-\hat{t}_{\text{MLE}})^{2}V_{A}+\frac{\hat{t}_{\text{MLE}}^{2}}{a^{2}}V_{A}+\hat{t}^{2}\hat{\xi}+1,$ (29)
@@
$a^{2}(\text{Var}(y^{\prime})-1)\approx\hat{t}_{\text{MLE}}^{2}V_{A}+a^{2}\hat{t^{2}}\hat{\xi}$, (30)
since the discrepancy $(\hat{t}-\hat{t}_{\text{MLE}})^{2}$ vanishes in the large-sample limit due to the consistency of the MLE. Thus, the covariance term $\text{Cov}(x,y^{\prime})$ is computed from the sample using the standard Pearson correlation estimator. The output of the network is then post-processed by dividing by $a^{2}$, recovering an accurate estimate of the original parameter $\hat{t}^{2}\hat{\xi}$ from the amplified noise features.
The use of an amplification factor $a>1$ to rescale input features is a well-established preprocessing technique in neural networks. However, its value must be chosen with physical insight to meaningfully amplify the parameter of interest while preserving numerical stability. In our context, $a$ scales the excess-noise contribution in the input feature $a^{2}(\text{Var}(y^{\prime})-1)$, making the excess noise more discernible against other signal components. The signal-to-noise ratio (SNR) of this amplified feature can be expressed as
$\mathrm{SNR}=10\log_{10}{\frac{TV_{\mathrm{A}}}{\mu+a^{2}T\xi}},$ (31)
For challenging regimes such as low transmittance ($T=0.1$) and high excess noise ($\xi=0.01$ SNU), the choice $a=10$ with $V_{A}=5\ \mathrm{SNU}$ yields $\mathrm{SNR}\approx-3.42$ dB, a value within a operational range of communication systems. Hence, $a=10$ represents a balanced selection that effectively elevates the excess-noise signal while maintaining estimation robustness across the considered parameter range.